OpenStudy (anonymous):

There are red and blue marbles in a jar. One-third of the marbles are red. If another marble is added to the jar, then 3/8 of the marbles will be red. How many marbles are blue?

OpenStudy (debbieg):

Set up the 2 proportions: \[\Large \frac{ R }{ N }=\frac{ 1 }{ 3 }\] and \[\Large \frac{ R }{ N+1 }=\frac{ 3 }{ 8 }\] And then solve the system with substitution.

OpenStudy (ybarrap):

$$ {r \over r+b}={1 \over3}\\ {r +1\over r+1+b}={3 \over8}\\ 3r=r+b \implies r={b\over 2}\\ 8(r+1)=3(r+1+b)\implies 5r=-5+3b\\ 5\left ({b \over 2 } \right )=-5+3b\\ \implies b=10 $$ That's it! Let me know if you have any questions.